OpenStudy (anonymous):

find the common ratio of each sequence. -3, -12, -48, -192

OpenStudy (raden):

ratio (r) = a2/a1 = a3/2 = a4/a3 = ..

OpenStudy (raden):

if u take r = a2/a1, here given the value of a2=-12 and a1=-3 so, r = a2/a1 = -12/-3 = 4

OpenStudy (raden):

it will same as the ration if u take : r = a3/a2 = -48/-12 = 4

OpenStudy (raden):

* ratio